OpenStudy (anonymous):

Solve: 4 over 5 y + 2 =6 5 7 9 11 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\large \frac{4}{5}y + 2 = 6\] right?

OpenStudy (anonymous):

5 (:

OpenStudy (anonymous):

Yes, @jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\large \frac{4}{5}y + 2 = 6\] \[\large \frac{4}{5}y = 6-2\] \[\large \frac{4}{5}y = 4\] \[\large 4y = 4*5\] \[\large 4y = 20\] \[\large y = \frac{20}{4}\] \[\large y = 5\]

OpenStudy (anonymous):

Thank you!